# Literature Review, Methodology, Software, Data curation, Implementation, Experimental analysis, Writing - Original draft preparation

### v1

@@ 3.2 Dataset splitting
For this experiment, we have split the dataset into 70-30% where 70% data is training data and 30% is testing data. In 70% training set total of 15,850 data exist and the testing set consists of 6,794 data. Initially, the training set was utilized for the train the model, and the testing test was used to find out the effectiveness of the model on unseen data identification and prediction is phishing or not.
@@ 3.3 Model Selection
In this experiment, we have employed the transformer-based BERT variant architecture DistilBERT which is a lightweight, fast, and economical model. This is developed for language processing tasks and also compresses the original BERT model and makes it faster. Knowledge distillation was utilized in the pre-training phase of this pre-trained model, which allowed it to be effectively reduced in size by 40% while maintaining 97% of its language understanding skills and delivering a 60% speed increase (Sanh et al., 2019). This small variant has twelve attention heads, 768 hidden layers, and six transformer layers (Adoma et al., 2020). Whereas DistilBERT contains about 66 million parameters, the original BERT model has 110M (million) parameters. By lowering the number of parameters, DistilBERT maintains excellent performance while achieving a smaller size and faster inference speed.
@@ 3.3.1 Model Optimization
To improve its handling of terms that are not in its lexicon and to capture meaningful sub-word representations, BERT uses a Word-Piece tokenizer, which splits words into smaller sub-words. The preprocessed phishing dataset is tokenized using the tokenizer of the DistilBERT transformer. In this tokenized portion, the maximum sequence length of input data has been used 512. Other crucial components of tokenization are padding and truncation. The major purpose of these procedures is to ensure that each tokenized input sequence has a constant length. This is important because it creates consistent embeddings and efficient processing in batches.
@@
Pre-trained weights acquired throughout the pre-training phase are used to initialize the DistilBERT models. Selecting an optimal batch size is paramount in deep learning, as it significantly influences both training and testing accuracies, along with overall runtime (Lin, 2022). The careful selection of batch size determines the efficiency of the neural network training process, striking a balance between computational resources and model performance. The batch sizes for training and testing data are optimized at 32 and 64, respectively. Each epoch’s training data is assigned at random to ensure the model sees a diversity of unseen data samples and prevents overfitting. This shuffling method increases the model’s ability to generalize.
@@
During training, an optimizer is essential in adjusting the parameters of a neural network. The main goal of this is to minimize the error or loss function of the model for better performance. For this research work, AdamW (Adam Weight Decay) (Loshchilov and Hutter, 2017) optimizer has been used to get better performance and avoid overfitting. This optimizer is a modification version of Adam which is widely adopted for training deep neural networks (Zhuang et al., 2022). By separating weight decay from the learning rate, AdamW enhances its implementation. It improves regularization and training dynamics by separating weight decay from gradient updates, in contrast to Adam. Learning rate is a crucial hyperparameter governing step size during training for optimizing models. The convergence and overall performance of the system are dependent on various aspects such as the model architecture, optimization techniques, and task domain. For efficient fine-tuning and fruitful model training, this parameter must be adjusted. Selecting a high learning rate can introduce instability, resulting in poor performance on unseen data, while a low learning rate may extend the training process, requiring more epochs and increasing computational costs. In this research, a learning rate, 2e-5 is used in the training process for the transformer-based DistilBERT model.
@@ 3.3.2 Fine Tuning
A pre-trained model is fine-tuned using task-specific data and backpropagation to alter parameters on certain tasks and datasets. The primary objective of this is to improve the model’s performance across various tasks and datasets while minimizing the loss function. In this experiment, DistilBERTForSequenceClassification has been utilized to adopt the pre-trained model. DataLoader is used to configure the train_loader and test_loader by setting prepared data, batch size, and shuffling command where train_loader is configured with shuffling and test_loader is configured without shuffling. After that, we fine-tuned this DistilBERT model by adjusting and adopting some necessary steps which are shown in Fig.  4.
Figure 4: The overall process of fine-tuning.
In this overall fine-tuning process, following the optimizer’s setup, the model is fine-tuned by feeding it the input parameters (input_ids, attention_mask, and label) during the forward pass so that predictions can be made. After the forward pass, gradients of the loss for the parameters (weights and biases) are calculated in the backward pass. These gradients show how changing the parameters would affect the loss. Lastly, an update is made to the model parameters. The optimizer is mainly responsible for changing the neural network’s parameters in accordance with the calculated gradients. It adjusts the weights and biases in an attempt to reduce the loss. The model gains the capacity to fine-tune its parameters throughout several epochs, which improves its predictive capability on the training set. The ultimate goal is to identify underlying patterns in the training set and use those patterns to generalize well to new data. Algorithm 1 shows the overall fine-tuning process in training.
Algorithm 1 Fine-Tuning the DistilBERT model
1:DistilBERT model, training data loader, optimizer, and number of epochs
2:Trained fine-tuned model
3:$\text{epochs}\Leftarrow n$
[…]
5:for epoch in range(epochsdo
6:     model.train()
7:     for batch in train_loader do
[…]
10:         $\text{labels}\Leftarrow\text{batch}[2]$
11:         $\text{gradients}\Leftarrow\text{0}$
[…]
13:         loss_calculation()
14:         backwardpass()
15:         update_parameters()
16:     end for
17:end for
@@ 3.3.3 Model Explainability
The ability of a model to explain its decisions is crucial since it provides insight into the process. Verifying the accuracy of judgments and reducing errors are made feasible by understanding how models work. Explainable AI has been a popular area with a growing group that has developed efficient techniques to understand deep neural network predictions among the more complex machine learning models (Holzinger et al., 2022). This is important because sophisticated models frequently act as opaque “black boxes" making it difficult for people to comprehend the logic behind particular decisions. In this research work, we have worked with model explanability to solve the black box problem. To understand the model prediction two explainability models LIME and Transformers Interpret have been utilized.
Generally, this experiment for phishing email detection is a text classification task that LimeTextExplainer used to explain the black box DistilBERT model. The purpose of the LIME method is to faithfully explain any classifier or regressor’s predictions (Ribeiro et al., 2016). For improved interpretability, LIME provides an understandable depiction of the decision-making process, providing light on the “why" behind the predictions or suggestions made by these models. In order to achieve model explainability, LIME first perturbs the email content by making modest adjustments, rearranging words, or adding new words. After that, the black-box transformer-based model is applied to the altered email content to generate predictions. From these impacted instances, LIME derives features. This local model should roughly represent the behaviour of the black-box model surrounding the chosen instance. Next, explanations for the black-box model’s prediction for the original data are produced using the local model. This explanation usually highlights the words or key elements most important during the model’s decision-making.
A Transformers Interpret tool was also utilized to explain the transformer-based model for this classification work. The class explainer process was carried out using this tool using a model and tokenizer. With the use of this class explanation, word attributions are computed. Positive attribution numbers signify a word’s positive contribution to the projected class, while negative numbers signify a word’s negative contribution. The terms that demonstrate a strong contribution to decision-making are also highlighted in this explanation. The overall process of model explainability is shown in Fig. 5. Algorithm 2 shows the explanation process of the model prediction.
@@
Algorithm 2 Explainability Analysis
1:Trained model, tokenizer, and text
@@
3:class_names $\Leftarrow$ [X, Y, …N]
@@
5:     LIME_Explanation(model, text)
6:     Transformer_Interpret(model, tokenizer, text)
7:end function
8:function LIME_Explanation(model, text)
9:     explainer $\Leftarrow$ LimeTextExplainer(class_names $\Leftarrow class\_names$)
10:     explanation $\Leftarrow$ explainer.explain_instance(text, model, num_features $\Leftarrow n$, top_labels $\Leftarrow 1$)
11:     exp_class $\Leftarrow$ explanation.available_labels()
12:     exp $\Leftarrow$ explanation.as_list($\Leftarrow$ exp_class)
13:     explanation_visualize()
14:end function
15:function Transformer_Interpret(model, tokenizer, text)
16:     cls_explainer $\Leftarrow$ SequenceClassificationExplainer(model, tokenizer)
17:     word_attributions $\Leftarrow$ cls_explainer(text)
18:     cls_explainer_visualize()
19:end function
@@ 4 Results
The fine-tuned transformer-based model was trained and tested with both unbalanced and balanced datasets. To assess the performance of the model different measurements such as confusion matrix, precision, recall, F1-score, and overall accuracy have been calculated. Overall summarize the performance, a model needs to calculate the confusion matrix first. It is a kind of table that provides a detailed breakdown of the actual and predicted classification for specific task domain data. The confusion matrix of the DistilBERT model for both imbalanced and balanced datasets is shown in Table 2. True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) are the main components of the confusion matrix and as well as are needed for the calculation of Precision, recall, and F1-score (Uddin et al., 2022). The precision, recall, and F1-score of the DistilBERT model for both imbalanced and balanced datasets are shown in Table 3 and Table 4.
Table 2: Confusion matrix of DistilBERT for both imbalanced and balanced dataset
@@
Class Name Safe Email Phishing Email Safe Email Phishing Email
Safe Email 3235 116 3281 98
Phishing Email 24 2216 5 3410
Table 3: The precision, recall, and f1-score for imbalanced dataset.
@@
Safe Email 0.99 0.97 0.98 3351
Phishing Email 0.95 0.99 0.97 2240
Table 4: The precision, recall, and f1-score for balanced dataset.
@@
Safe Email 1.00 0.97 0.98 3379
Phishing Email 0.97 1.00 0.99 3415
Precision: Precision which expresses the ratio of true positives to all predicted positives, indicates how accurate positive predictions are and emphasizes the model’s capacity to identify positive samples.
@@
F1-score: A balanced statistical metric that assesses a model’s efficaciousness in predicting positive instances while accounting for imbalanced datasets is the F1-score, which is derived from the harmonic mean of accuracy and recall. It offers a thorough perspective that balances recall and precision, impacting the final score determined by the model’s projections of the positive rate and actual positive cases.
@@
The accuracy of a model can differ based on the various parameters, data balanced or not. Initially in this experiment, an imbalanced dataset was utilized to assess the model’s performance in both training and testing. This dataset is highly imbalanced based on the safe email and phishing email data where the number of data of safe email is quite higher than phishing email. To train the model with 70% imbalanced data we have used 6 epochs with a batch size of 32. After 6 epochs the model showed 98.90% training accuracy for this imbalanced dataset with 0.0201 training loss. On the other hand, we get a testing accuracy 97.50 with testing loss 0.0900 with batch size 64. Fig.  6 shows a graph of accuracy and loss with 6 epochs for an imbalanced dataset.
Figure 6: Training vs. Testing Accuracy and Loss on Imbalanced Data
Moreover, we also conducted this experiment with a resampled and balanced dataset. After preparing the balanced dataset again used 70% of the data for training purposes. Using this portion of data, we worked on model optimization and fine-tuning and trained the model. This time with balanced data the fine-tuned model showed 99.07% training accuracy with 0.0218 training loss. After that, we tested this trained model with 30% unseen balanced data and got 98.48% testing accuracy with 0.0447 testing loss. This performance showed that the performance of the model with a balanced dataset is quite better than an imbalanced dataset. Fig. 7 shows a graph of accuracy and loss with 6 epochs for a balanced dataset. The performance comparison of the model on both balanced and imbalanced dataset shows in Table 5.
Figure 7: Training vs. Testing Accuracy and Loss on Imbalanced Data.
Table 5: The Performance comparison of fine-tuned DistilBERT model on Imbalanced and Balanced dataset.
@@
Training Loss
Training Accuracy
Testing Loss
Testing Accuracy
Imbalanced
0.0201
98.90
0.0900
97.50
Balanced
0.0218
99.07
0.0447
98.48
@@
Reference
Datasets
Models
Explainability Analysis
2023(Jamal et al., 2023)
Email Spam Detection Dataset, PhishingEmailData
DistilBERT, RoBERTA
No
2021(Alhogail and Alsabih, 2021)
Fraud Dataset
GCN
No
2019(Fang et al., 2019)
IWSPA-AP
THEMIS
No
2023(Thapa et al., 2023)
IWSPA-AP, Nazario, Enron, CSIRO, Phisbowl
THEMIS, BERT
No
2023(Atawneh and Aljehani, 2023)
Phishing Email Dataset, Benign Email Dataset
BERT-LSTM
No
This Study Phishing Email Fine-tuned DistilBERT Yes
The comparison showed that our research work provides transparency in this phishing email detection task. However, some listed research work showed higher accuracy than our proposed model. This happened because they used different datasets in their experiment. In this study, we only focused on the phishing email dataset but other existing works were also done this work with the spam dataset. Furthermore, to the best of our knowledge, none of the papers listed in Table  6 offer a thorough human explanation of the classifier’s performance or clarify how the model functions in phishing email detection.

### v2

@@ 3.2 Dataset splitting
For this experiment, we split the dataset into 80-20%, where 80% of the data is used for training and 20% for testing. The 80% training data is then further split into 80-20%, resulting in an overall split of 64% training set, 16% validation set, and 20% testing set. The 64% training set contains a total of 14,492 data, the validation set consists of 3,623 data, and the testing set comprises 4,529 data. Initially, the training set was used to train the model, the validation set was used to validate the model to ensure its generalizability, and the testing set was used to evaluate the model’s effectiveness in identifying and predicting phishing on unseen data.
@@ 3.3 Model Selection
In this experiment, we have employed RoBERTa, a transformer-based robustly optimized variant of BERT, which retains BERT’s architecture and has refined the training process through key modifications. RoBERTa is trained on a much larger text dataset (160 GB of text) using 500,000 training steps with a large batch size of 8,000 per unit(Liu et al., 2019). The training process benefited greatly from dynamic masking, which updates the masking scheme during training instead of using static masking like BERT did during data preprocessing. This allows the model to learn a much more diverse set of training examples (from the same text) and also improves robustness. Moreover, the model removes the Next Sentence Prediction (NSP) task that BERT used, only uses MLM, finding it unnecessary and potentially less helpful to performance. Additionally, it uses a more efficient tokenization method, Byte-Level Byte Pair Encoding (BPE), similar to GPT-2, which helps handle rare words better. The RoBERTa-base contains about 125 million parameters. This variant has 12 attention heads, a 768-dimensional hidden state, and 12 transformer layers. The architecture of the RoBERTa model is illustrated in Fig.  4.
Figure 4: The architecture of the RoBERTa model
@@ 3.3.1 Model Optimization
RoBERTa employs a BPE tokenizer that is capable of splitting words into smaller sub-word units to better manage out-of-vocabulary terms and to allow for more meaningful sub-word representations. The preprocessed phishing dataset is tokenized using RoBERTaTokenizer, the default tokenizer for the RoBERTa transformer model. The maximum sequence length of the input data is configured to be a maximum of 512 tokens during the tokenization process, as this is the standard length for most transformer-based architectures, including RoBERTa. Moreover, as we are working with an email dataset, where the email texts are often long, setting the token length to less than 512 could result in a potential loss of important information. Other important steps in the tokenization process include padding and truncation: padding fills in sequences that are shorter than the maximum length with special token padding tokens, and truncation removes tokens from the end of sequences that are longer than the maximum length. These two tasks are important as they ensure input dimensions are standardized across the algorithm, so that embeddings are consistent and the batches of data can be processed efficiently during training and inference time.
@@
Pre-trained weights acquired throughout the pre-training phase are used to initialize the RoBERTa model. Selecting an optimal batch size is paramount in deep learning, as it significantly influences both training and testing accuracies, along with overall runtime (Lin, 2022). The careful selection of batch size determines the efficiency of the neural network training process, striking a balance between computational resources and model performance. The batch sizes for training and testing data are optimized at 32 and 64, respectively. Each epoch’s training data is assigned at random to ensure the model sees a diversity of unseen data samples and prevents overfitting. This shuffling method increases the model’s ability to generalize.
@@
During training, an optimizer is essential in adjusting the parameters of a neural network. The main goal of this is to minimize the error or loss function of the model for better performance. For this research work, AdamW (Adam Weight Decay) (Loshchilov and Hutter, 2017) optimizer has been used to get better performance and avoid overfitting. This optimizer is a modified version of Adam, which is widely adopted for training deep neural networks (Zhuang et al., 2022). By separating weight decay from the learning rate, AdamW enhances its implementation. It improves regularization and training dynamics by separating weight decay from gradient updates, in contrast to Adam. Learning rate is a crucial hyperparameter governing the step size during training for optimizing models. The convergence and overall performance of the system are dependent on various aspects such as the model architecture, optimization techniques, and task domain. For efficient fine-tuning and fruitful model training, this parameter must be adjusted. Selecting a high learning rate can introduce instability, resulting in poor performance on unseen data, while a low learning rate may extend the training process, requiring more epochs and increasing computational costs. In this research, a learning rate, 2e-5 is used in the training process for the transformer-based RoBERTa model.
@@ 3.3.2 Fine Tuning
A pre-trained model is fine-tuned using task-specific data and backpropagation to alter parameters on certain tasks and datasets. The primary objective of this is to improve the model’s performance across various tasks and datasets while minimizing the loss function. In this experiment, RobertaForSequenceClassification has been utilized to adopt the pre-trained model. DataLoader is used to configure the train_loader and test_loader by setting prepared data, batch size, and shuffling command, where train_loader is configured with shuffling and test_loader is configured without shuffling. After that, we fine-tuned this RoBERTa model during the training process by adjusting and adopting some necessary steps: optimizer setup, forward pass, loss calculation, backward pass (backpropagation), and updating model parameters, which are shown in Fig. 5.
Figure 5: The process of fine-tuning.
In this overall fine-tuning process, following the optimizer’s setup, the model is fine-tuned by feeding it the input parameters (input_ids, attention_mask, and label) during the forward pass so that predictions can be made. After the forward pass, gradients of the loss for the parameters (weights and biases) are calculated in the backward pass. These gradients show how changing the parameters would affect the loss. Lastly, an update is made to the model parameters. The optimizer is mainly responsible for changing the neural network’s parameters in accordance with the calculated gradients. It adjusts the weights and biases in an attempt to reduce the loss. The model gains the capacity to fine-tune its parameters throughout several epochs, which improves its predictive capability on the training set. The ultimate goal is to identify underlying patterns in the training set and use those patterns to generalize well to new data.
@@ 3.3.3 Model Explainability
The ability of a model to explain its decisions is crucial since it provides insight into the process. Verifying the accuracy of judgments and reducing errors are made feasible by understanding how models work. XAI has been a popular area with a growing group that has developed efficient techniques to understand deep neural network predictions among the more complex machine learning models (Holzinger et al., 2022). This is important because sophisticated models frequently act as opaque “black boxes" making it difficult for people to comprehend the logic behind particular decisions Sarker (2024). In this research work, we have worked with model explainability to solve the black box problem. To understand the model prediction, two popular explainability models, LIME and Transformers Interpret, have been utilized. The overall process of model explainability is shown in Fig. 6.
@@
Generally, this experiment for phishing email detection is a text classification task that LimeTextExplainer used to explain the black box RoBERTa model. The purpose of the LIME method is to faithfully explain any classifier or regressor’s predictions (Ribeiro et al., 2016). For improved interpretability, LIME provides an understandable depiction of the decision-making process, providing light on the “why" behind the predictions or suggestions made by these models. To achieve model explainability, LIME first perturbs the email content by making modest adjustments, such as rearranging words or adding new ones. After that, the black-box transformer-based model is applied to the altered email content to generate predictions. From these impacted instances, LIME derives features. This local model should roughly represent the behaviour of the black-box model surrounding the chosen instance. Next, explanations for the black-box model’s prediction for the original data are produced using the local model. This explanation usually highlights the words or key elements most important during the model’s decision-making process.
A Transformers Interpret tool was also utilized to explain the transformer-based model for this classification work. The class explainer process was carried out using this tool, using a model and tokenizer. With the use of this class explanation, word attributions are computed. Positive attribution numbers signify a word’s positive contribution to the projected class, while negative numbers signify a word’s negative contribution. The terms that demonstrate a strong contribution to decision-making are also highlighted in this explanation.
However, while LIME and Transformers Interpret both provide a suitable understanding of a model’s decision-making processes, they differ fundamentally in their methods for explanation. LIME operates on word-level perturbations, whereas Transformers Interpret uses gradient-based attribution on subword tokens derived from the tokenizer. As a result, the outputs of these techniques, such as the number of highlighted tokens, the magnitude of scores, and their interpretive granularity, often vary, potentially leading to confusion for users. To solve this limitation and enhance the clarity and consistency of the explanations, we present a hybrid explanation approach called LITA. This method integrates the attribution scores from both LIME and Transformers Interpret into a unified explanation by computing a weighted combination of their normalized outputs. Algorithm  1 shows the explanation process of the model prediction using a hybrid approach to generate comprehensive explanatory scores. Here, $\alpha$ is a tunable parameter in the range $[0,1]$, controlling the influence of each technique. Both LIME (norm_lime_scores) and Transformers (norm_transformer_scores) attribution scores are first normalized to a common range [1,1][-1,1] to ensure consistency and comparability. As we are not sure which method is more accurate so we set $\alpha=0.5$ in this experiment, which assigns equal weight to both methods. In addition to reducing the overall differences between both attribution techniques, offering an indication of total feature importance with a hybrid attribution technique offers an additional layer of interpretability and reliability of the decision-making process.
Algorithm 1 Hybrid Explainability Method
1:Trained model, tokenizer, and input text
@@
3:class_names $\Leftarrow$ [X, Y, …, N]
@@
5:  lime_scores $\Leftarrow$ LIME_Explanation(model, text)
6:  transformer_scores $\Leftarrow$ Transformer_Interpret(model, tokenizer, text)
7:  normalized_lime $\Leftarrow$ lime_scores scaled to [-1, 1]
8:  normalized_transformer $\Leftarrow$ transformer_scores scaled to [-1, 1]
9:  hybrid_scores $\Leftarrow$ Hybrid_Explanation(normalized_lime, normalized_transformer, $\alpha$)
10:  return lime_scores, transformer_scores, hybrid_scores
11:end function
12:function LIME_Explanation(model, text)
13:  explainer $\Leftarrow$ LimeTextExplainer(class_names $\Leftarrow$ class_names)
14:  explanation $\Leftarrow$ explainer.explain_instance(text, model, num_features $\Leftarrow n$, top_labels $\Leftarrow 1$)
15:  exp_class $\Leftarrow$ explanation.available_labels()
16:  lime_scores $\Leftarrow$ explanation.as_list(label $\Leftarrow$ exp_class)
17:  return lime_scores
18:end function
19:function Transformer_Interpret(model, tokenizer, text)
20:  cls_explainer $\Leftarrow$ SequenceClassificationExplainer(model, tokenizer)
21:  transformer_scores $\Leftarrow$ cls_explainer(text)
22:  return transformer_scores
23:end function
24:function Hybrid_Explanation(norm_lime_scores, norm_transformer_scores, $\alpha$)
25:  for all word in norm_lime_scores $\cap$ norm_transformer_scores do
26:   hybrid[word] $\Leftarrow\alpha\cdot$ norm_lime_scores[word] $(1-\alpha)\cdot$ norm_transformer_scores[word]
27:  end for
28:  return hybrid
29:end function
@@ 4 Results
A model’s generalization power can be assessed using various performance metrics, including precision, recall, F1-score, and accuracy, across different conditions (Chen et al., 2024). The fine-tuned transformer-based model was trained and tested with unbalanced and balanced datasets. To assess the performance of the model, different measurements such as the confusion matrix, precision, recall, F1-score, and overall accuracy have been calculated. Overall summarize the performance, a model needs to calculate the confusion matrix first. This kind of table provides a detailed breakdown of the actual and predicted classification for specific task domain data. The RoBERTa model’s confusion matrix for imbalanced and balanced datasets is shown in Table 2. True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) are the main components of the confusion matrix, and as well as are also needed for the calculation of Precision, recall, and F1-score (Uddin et al., 2022). The precision, recall, and F1-score of the RoBERTa model for both imbalanced and balanced datasets are shown in Table 3 and Table 4.
Table 2: Confusion matrix of RoBERTa for both imbalanced and balanced datasets
@@
Class Name Legitimate Email Phishing Email Legitimate Email Phishing Email
Legitimate Email 2150 59 2181 60
Phishing Email 17 1501 10 2278
Table 3: The precision, recall, and f1-score for RoBERTa model on imbalanced dataset.
@@
Legitimate Email 0.99 0.97 0.98 2209
Phishing Email 0.96 0.99 0.98 1518
Table 4: The precision, recall, and f1-score for RoBERTa model on balanced dataset.
@@
Legitimate Email 1.00 0.97 0.98 2241
Phishing Email 0.97 1.00 0.98 2288
Precision: Precision, which expresses the ratio of true positives to all predicted positives, indicates how accurate positive predictions are and emphasizes the model’s capacity to identify positive samples.
@@
F1-score: A balanced statistical metric that assesses a model’s efficacy in predicting positive instances while accounting for imbalanced datasets is the F1-score, which is derived from the harmonic mean of accuracy and recall. It offers a thorough perspective that balances recall and precision, impacting the final score determined by the model’s projections of the positive rate and actual positive cases.
@@
The accuracy of a model can differ based on the various parameters, data is balanced or not. Initially, in this experiment, an imbalanced dataset was utilized to assess the model’s performance in training, validation, and testing. This dataset is highly imbalanced based on the legitimate email and phishing email data, where the number of data of legitimate emails is much higher than phishing emails. To train the model with 64% imbalanced data, we have used 5 epochs with a batch size of 32. After 5 epochs, the model showed 98.57% training accuracy for this imbalanced dataset with 0.0281 training loss. Then the trained model was validated with 16% unseen data and achieved 98.05% validation accuracy with 0.0513 loss. On the other hand, we got a testing accuracy 97.96 with testing loss 0.0560 with batch size 64. Fig.  7 shows a graph of accuracy and loss with 6 epochs for an imbalanced dataset.
Figure 7: Training vs. Validation vs. Testing Accuracy and Loss on Imbalanced Data
Moreover, we also conducted this experiment with a balanced dataset. After preparing the balanced dataset again used 64% of the data for training purposes. We worked on model optimization and fine-tuning using this data portion and trained the model. With the balanced data, the fine-tuned model achieved 98.94% training accuracy with 0.0257 training loss, indicating exceptional performance on the training data. After that, the fine-tuned model was validated with 16% validation data and obtained 0.0388 validation loss with 98.68% validation accuracy, demonstrating that the model generalizes well to unseen data and is not overfitting. Finally, the trained model was tested with 20% unseen balanced test data and achieved 98.45% testing accuracy with 0.0511 testing loss. This performance showed that the performance of the model with a balanced dataset is much better than an imbalanced dataset. Fig. 8 shows a graph of accuracy and loss with 5 epochs for a balanced dataset. The performance comparison of the model on both balanced and imbalanced datasets is shown in Table 5.
Figure 8: Training vs. Validation vs. Testing Accuracy and Loss on Balanced Data.
Table 5: The Performance Comparison of Fine-Tuned RoBERTa Model on Imbalanced and Balanced Datasets.
@@
Imbalanced Balanced
Training Loss 0.0281 0.0257
Training Accuracy 98.57 98.94
Validation Loss 0.0513 0.0388
Validation Accuracy 98.05 98.68
Testing Loss 0.0560 0.0511
Testing Accuracy 97.96 98.45
Additionally, we employed traditional machine learning models, including Extreme Gradient Boosting (XGB), Random Forest (RF), Support Vector Machine (SVM), and Logistic Regression (LR), to evaluate their performance and compare them with our fine-tuned RoBERTa model on an imbalanced and balanced dataset. The generalization power and robustness of the different ML models and the RoBERTa model are shown in Table 6 and Table 7. Table 8 and Table 9 present a comparative performance analysis of these models. The overall performance analysis showed that our fine-tuned RoBERTa model outperformed the traditional ML models.
Table 6: Precision, Recall, and F1-Score of traditional ML and RoBERTa Model for Imbalanced Dataset
Model Class Precision Recall F1-score
RF Legitimate Email 0.96 0.97 0.96
Phishing Email 0.95 0.94 0.95
XGB Legitimate Email 0.98 0.95 0.96
Phishing Email 0.93 0.98 0.95
SVM Legitimate Email 0.99 0.97 0.98
Phishing Email 0.95 0.99 0.97
LR Legitimate Email 0.98 0.96 0.97
Phishing Email 0.94 0.97 0.96
RoBERTa Legitimate Email 0.99 0.97 0.98
Phishing Email 0.96 0.99 0.98
Table 7: Precision, Recall, and F1-Score of traditional ML and RoBERTa Model for Balanced Dataset
Model Class Precision Recall F1-score
RF Legitimate Email 0.98 0.96 0.97
Phishing Email 0.96 0.98 0.97
XGB Legitimate Email 0.99 0.94 0.96
Phishing Email 0.95 0.99 0.97
SVM Legitimate Email 0.99 0.97 0.98
Phishing Email 0.97 0.99 0.98
LR Legitimate Email 0.99 0.96 0.97
Phishing Email 0.96 0.99 0.97
RoBERTa Legitimate Email 1.00 0.97 0.98
Phishing Email 0.97 1.00 0.98
Table 8: Comparative analysis of the fine-tuned RoBERTa model with traditional ML models on the imbalanced dataset.
Model Training Loss Training Accuracy Testing Loss Testing Accuracy
RF 0.0834 98.91 0.2241 95.65
XGB 0.0684 98.02 0.1071 95.84
SVM 0.0531 98.77 0.0864 97.61
LR 0.1446 97.89 0.1707 96.46
RoBERTa 0.0281 98.57 0.0560 97.96
Table 9: Comparative analysis of the fine-tuned RoBERTa model with traditional ML models on the balanced dataset.
Model Training Loss Training Accuracy Testing Loss Testing Accuracy
RF 0.0726 99.17 0.1837 97.09
XGB 0.0666 98.32 0.0966 96.49
SVM 0.0422 99.01 0.0710 97.90
LR 0.1343 98.36 0.1489 97.24
RoBERTa 0.0257 98.94 0.0511 98.45
@@
Reference Datasets Models Explainability Analysis
2023(Jamal et al., 2023) Email Spam Detection Dataset, PhishingEmailData DistilBERT, RoBERTA No
2021(Alhogail and Alsabih, 2021) Fraud Dataset GCN No
2019(Fang et al., 2019) IWSPA-AP THEMIS No
2023(Thapa et al., 2023) IWSPA-AP, Nazario, Enron, CSIRO, Phisbowl THEMIS, BERT No
2023(Atawneh and Aljehani, 2023) Phishing Email Dataset, Benign Email Dataset BERT-LSTM No
This Study Phishing Email Fine-tuned RoBERTa Yes (LIME, Transformers Interpret, Hybrid LITA)
The comparison showed that our research work provides transparency in the phishing email detection task. However, some of the listed studies showed higher accuracy than our proposed model. This happened because they used different datasets in their experiment. In this study, we only focused on the phishing email dataset, whereas other existing works also included the spam datasets. Furthermore, to the best of our knowledge, none of the papers listed in Table  10 offer a thorough human explanation of the classifier’s performance or clarify how the model functions in phishing email detection.